OpenStudy (anonymous):

Solve: (tan x - 1)(tan x - √3) = 0 0≤Ɵ<2π

OpenStudy (anonymous):

tanx-1=0 tanx-sqrt(3)=0 tanx=1 tanx=sqrt(3) for tanx=1, x=pi/4 tangent is also positive in the 3rd quadrant, so subtract pi/4 from 3pi/2 to get your second answer for this half of the equation. follow the same rules for tanx=sqrt(3)

OpenStudy (anonymous):

Where is the 3pi/2 come from ?

OpenStudy (anonymous):

3pi/2-pi/4 will give you a solution in the 3rd quadrant

OpenStudy (anonymous):

where tangent is also positive

OpenStudy (anonymous):

its the same as pi/4+pi

OpenStudy (anonymous):

pi/4+pi=3pi/2-pi/4=5pi/4

OpenStudy (anonymous):

Sorry if it's a silly questions but how is pi/4 + pi = 3pi/2 ? can you please explain this part, cuz I'm really confuse in class too about all this pi thing

OpenStudy (anonymous):

\[\frac{\pi}{4}+\pi=\frac{\pi}{4}+\frac{4\pi}{4}=\frac{\pi+4\pi}{4}=\frac{5\pi}{4}\]

OpenStudy (anonymous):

pi/4+pi does not equal 3pi/2 it equals 3pi/2-pi/4

OpenStudy (anonymous):

both pi/4+pi and 3pi/2-pi/4 equal 5pi/4

OpenStudy (anonymous):

see how i multiplied pi above by 4/4, this is to get each term with an LCD

OpenStudy (anonymous):

then the addition is simple

OpenStudy (anonymous):

i've only given you 2 solutions so far. pi/4 and 5pi/4. there are 2 more solutions resulting from: tanx=sqrt(3)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i got one it's pi/3

OpenStudy (anonymous):

yes, and the 2nd one is?

OpenStudy (anonymous):

add pi to pi/3

OpenStudy (anonymous):

4pi/3

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks you so much Dockworker ! :D

OpenStudy (anonymous):

those are all solutions in the domain 0<=x<=2pi

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

yw